OpenStudy (anonymous):

Write an expression for the volume of a cylinder with a height 7in. greater than the radius.

OpenStudy (tkhunny):

You need somewhere to start. How about the Radius measures "r". This makes the height "r + 7"

OpenStudy (anonymous):

Radius= r Height= h So: h= r+7

OpenStudy (anonymous):

Volume of Cylinder= pi × r^2 × h

OpenStudy (anonymous):

r= h-7

OpenStudy (anonymous):

Value of height hasn't be provided so the answer remains as an expression: pi × (h-7)^2 × h